OpenStudy (anonymous):

Rewrite using only sine and cosine: sin 3x I know the double angle formula and how to use it, could someone walk me through it though so I understand it completely?

OpenStudy (anonymous):

\[\sin(3x)= \sin (2x + x)\] remember that \[\sin (A +B) = sinAcosB + sinBcosA\]

OpenStudy (anonymous):

So in this case \[\sin(3x) = \sin(2x)\cos(x) + \sin(x)\cos(2x)\] which is \[2\sin(x)\cos ^{2}x + \sin(x)(1-2\sin^{2}(x))\]

OpenStudy (anonymous):

do you follow how I used the sin2x = 2sinxcosx and the cos2x = 1- 2sin^2(x) ?

OpenStudy (anonymous):

So you broke sin(3x) to create two double angle formulas of sin(2x) and cos(2x)?

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I broke it into one double angle formula ... the other one was just a result of breaking the ladder...